OpenStudy (anonymous):

How will adding the value 2 affect the mean and median of the data set 4, 4, 4, 6, 8? A. The mean and median will decrease by the same amount. B. The median will decrease more than the mean. C. The median will stay the same, but the mean will decrease. D. The mean and median will both stay the same.

OpenStudy (anonymous):

@jtbulin

OpenStudy (anonymous):

@72595210002 lol

OpenStudy (anonymous):

what

OpenStudy (anonymous):

@72595210002 why don't you try to solve it?

OpenStudy (anonymous):

alright, mean= 4+4+4+6+8= 26 That is 26/5= 5.2 .. if you add 2, it is 28/6=4.6..the difference is .6

OpenStudy (anonymous):

it is C

OpenStudy (anonymous):

now lets do the median its 4. now add the 2 its still 4..

OpenStudy (anonymous):

you are right, it is C